A stock just paid annual dividends of $3.55 per share. The dividends are expected to grow at 20 per cent per year for 4 years, and then remain constant in perpetuity. If the investors' required return for the stock is 11.8 percent, what should be the price of the stock today?
Answer #1

D1 = 3.55(1.20) = $4.26

D2 = 3.55(1.20)2 = $5.11

D3 = 3.55(1.20)3 = $6.13

D4 = 3.55(1.20)4 = $7.36

Stock Price = 4.26/(1.118) + 5.11/(1.118)2 + 6.13/(1.118)3 + 7.36/(1.118)4 + 7.36/(0.118)(1.118)4

Stock Price = $56.92